OpenStudy (anonymous):

What is ln0? What is ln0? @Mathematics

OpenStudy (anonymous):

undefined since For all real number of x, 10^x can not be = 0

OpenStudy (anonymous):

So I agree with moneybird^

jimthompson5910 (jim_thompson5910):

yes, ln(0) is undefined since ln(0) = y implies e^y = 0, but there are no values of y which satisfy this equation

OpenStudy (turingtest):

yeah or \[-\infty\]

jimthompson5910 (jim_thompson5910):

in a calculus context, you can say that \[\Large \lim_{x\to0+}\left(\ln(x)\right)=-\infty\] but not sure the book is looking for that

OpenStudy (turingtest):

Sure, sometimes they want "undefined" and sometimes they want "infinity". It can be hard to tell what they are looking for without more context.

OpenStudy (anonymous):

Before calculus you are supposed to say it's undefined as you don't know what values it is going to. After learning limits, most teachers want you to say Infinity

OpenStudy (anonymous):

depends on the level of your class